During the past year, the high and low use of three different resources for Fly High Airlines occurred in July and April. The resources are airplane depreciation, fuel, and airplane maintenance. The number of airplane flight hours is the driver. The total costs of the three resources and the related number of airplane flight hours are as follows:


Resource   Airplane Flight
Hours  
Total Cost
Airplane depreciation:                 
High   44,000         $ 18,700,000     
Low   28,000         $ 18,700,000     
Fuel:                 
High   44,000         445,896,000     
Low   28,000         283,752,000     
Airplane maintenance:                 
High   44,000         15,864,000     
Low   28,000         11,912,000     
Required:

Use the high-low method to answer the following questions. If an answer is zero, enter "0". If required, round your answers to the nearest dollar.

1. What is the variable-rate for airplane depreciation?

per flight hour

What is the fixed cost for airplane depreciation?

2. What is the cost formula for airplane depreciation?

Total cost of airplane depreciation =


3. What is the variable-rate for fuel?
$

What is the fixed cost for fuel?

4. What is the cost formula for fuel?

Total cost of fuel = $ x Number of airplane flight hours

5. What is the variable-rate for airplane maintenance?
$
per flight hour

What is the fixed cost of airplane maintenance?
$
6. What is the cost formula for airplane maintenance?

7. Using the three cost formulas that you developed, predict the cost of each resource in a month with 36,000 airplane flight hours. (Note: Do not round intermediate calculations.)

Total cost of airplane depreciation   $
Total cost of fuel   $
Total cost of airplane maintenance   $
Answer #1

1. Airplane depreciation:

  Variable rate

= $18,700,000 – $18,700,000. /  44000 – 28000= $0

  Fixed cost = $18,700,000 – ($0 × 44,000) = $18,700,000

2. Total cost of airplane depreciation = $18,700,000

  Airplane depreciation is a strictly fixed cost.

3. Fuel:

  Variable rate = $445,896,000 - $283,752,000 44,000 - 28,000 = $10,134

  Fixed cost = $445,896,000 – ($10,134 × 44,000) = $0

4. Total cost of fuel = $10,134 × Number of airplane flight hours

  Fuel is a strictly variable cost.

5. Airplane maintenance:

  Variable rate

=$15,864,000 – $11,912,000. / 44000 – 28000   = $247

  Fixed cost = $15,864,000 – ($247 × 44,000) = $4,996,000.

6. Total cost of airplane maintenance = $4,996,000 + ($247 × Number of airplane flight hours)

  Airplane maintenance is a mixed cost.

7. Total cost of each resource at 36,000 machine hours:

  Total cost of airplane depreciation = $18,700,000

  Total cost of fuel = $10,134(36,000) = $364,824,000

  Total cost of airplane maintenance = $4,996,000 + ($247 × 36,000) = $13,888,000.